OpenStudy (anonymous):

solve 1/2=(b/4)^(1/4)

OpenStudy (anonymous):

b= 1/4

OpenStudy (anonymous):

wouldnt it just be: \[1/2^{4}=b/4\]

OpenStudy (anonymous):

b=1/4

OpenStudy (anonymous):

yep- so you can just flip the exponent over when you go to the other side of the equation?

OpenStudy (anonymous):

indeed, 1/2^4 = b/4 means 2^4*b=4 so b=1/4

OpenStudy (anonymous):

awesome, thank you